### v1

# Class-Based Expurgation Attains Csiszár’s Expurgated Source-Channel Exponent

Abstract
This paper studies expurgated error exponents for joint source-channel coding for discrete memoryless sources and channels. We consider a partition of the source messages into classes, where the codeword distributions depend on the class. We show that two carefully chosen classes suffice to achieve Csiszár’s expurgated exponent.
I Introduction
We study the transmission of non-equiprobable messages from a discrete memoryless source (DMS) with distribution $P^{k}(\boldsymbol{v})=\prod_{i=1}^{k}P_{V}(v_{i})$, where $\boldsymbol{v}=(v_{1},...,v_{k})\in\mathcal{V}^{k}$ is the source message, and $\mathcal{V}$ is a finite discrete alphabet. The channel is a discrete memoryless channel (DMC) given by $W^{n}(\boldsymbol{y}|\boldsymbol{x})=\prod_{i=1}^{n}W(y_{i}|x_{i})$, $\boldsymbol{x}=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$ and $\boldsymbol{y}=(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}$, where $\mathcal{X}$ and $\mathcal{Y}$ are discrete alphabets with cardinalities $|\mathcal{X}|$ and $|\mathcal{Y}|$, respectively. An encoder maps the length-$k$ source message $\boldsymbol{v}$ to a length-$n$ codeword $\boldsymbol{x}_{\boldsymbol{v}}$, which is then transmitted over the channel. We refer to $t\triangleq k/n$ as the transmission rate. Based on the channel output $\boldsymbol{y}$ the decoder guesses which source message was transmitted.
We say that an error exponent $E>0$ is achievable if there exists a sequence of codes of length $n$ such that the error probability satisfies
$p_{e}\leq e^{-nE+o(n)},$ (1)
where $\lim_{n\rightarrow\infty}o(n)/n=0$. Most achievability results are obtained using random coding, followed by analyzing the average probability of error. It has been shown that joint source-channel coding (JSCC) can achieve lower error probabilities than separate source and channel coding [1]. In Csiszár’s construction, codewords are randomly selected from a set of sequences whose composition depends on the source message composition. Later, Zhong et al.[2] studied the random coding error exponent in JSCC in more detail and proved that Csiszár’s exponent corresponds to the concave hull of the $E_{0}$ function in the dual domain. A more general message set partitioning method was studied in [3], where it was shown that two carefully chosen classes are enough to achieve Csiszár’s random coding exponent.
Expurgation has received less attention than random coding bounds in JSCC. In [4], Csiszár derived two expurgated exponents, which were later recovered by Scarlett [5] using Gallager’s classical expurgation method [6, Sec. 5.7]. Motivated by the preceding discussion, we study expurgation under a partition of the source messages. We show that, as in the random-coding case, two carefully chosen classes are sufficient to achieve Csiszár’s second expurgated exponent [4].
Notation: In this paper, scalar random variables are denoted by uppercase letters, their realizations by lowercase letters, and their alphabets by calligraphic letters. Random vectors are written in boldface. For two positive sequences $f_{n}$ and $g_{n}$, we write $f_{n}\doteq g_{n}$ if $\lim_{n\rightarrow\infty}\frac{1}{n}\log\frac{f_{n}}{g_{n}}=0$, and we write $f_{n}\,\dot{\leq}\,\,g_{n}$ if $\lim\sup_{n\rightarrow\infty}\frac{1}{n}\log\frac{f_{n}}{g_{n}}\leq 0$.
The type of a sequence $\boldsymbol{x}=\left(x_{1},...,x_{n}\right)\in\mathcal{X}^{n}$ is its empirical distribution, defined by $\hat{P}_{\boldsymbol{x}}(x)\triangleq\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\left\{{x}_{i}=x\right\}.$ The set of all probability distributions on an alphabet $\mathcal{X}$ is denoted by $\mathcal{P}(\mathcal{X})$, while $\mathcal{P}_{n}(\mathcal{X})$ represents the set of empirical distributions for vectors in $\mathcal{X}^{n}$. For $P_{X}\in\mathcal{P}_{n}(\mathcal{X})$, the type class $\mathcal{T}^{n}(P_{X})$ consists of all sequences in $\mathcal{X}^{n}$ with type $P_{X}$.
The number of source type classes in the set $\mathcal{V}^{k}$ is denoted by $N_{k}$, which grows polynomially with $k$ [7, Lemma 2.2]. Throughout, we use the indices $i$ and $j$ to refer to source types, and $c$ to indicate the class to which a source message belongs under a given partitioning. The total number of classes in the partitioning is denoted by $m$. Moreover, for a source type $P_{i}$, we define $R_{i}\triangleq tH(P_{i})$.
II Csiszár’s Expurgated Error Exponents
In this section, we introduce definitions and summarize existing expurgated results in JSCC that are relevant to this work. The source reliability function [7, Ch. 9] is given by
$e(R,P_{V})$ $\triangleq\min_{Q:H(Q)\geq R}D(Q\|P_{V}),$ (2)
or, in the dual domain [8], as
$e(R,P_{V})$ $=\sup_{\rho\geq 0}\Big{\{}\rho R-E_{s}(\rho,P_{V})\Big{\}},$ (3)
where
$E_{s}(\rho,P_{V})$ $\triangleq\log\bigg{(}\sum_{v\in\mathcal{V}}P_{V}(v)^{\frac{1}{1+\rho}}\bigg{)}^{1+\rho}.$ (4)
The weak channel expurgated exponent is given by
$\!E_{\mathrm{ex}}^{\prime}(Q,R)\!\triangleq\!\!\min_{\begin{subarray}{c}P_{X\bar{X}}:P_{X}=Q\\ I_{P}(X;\bar{X})\leq R\end{subarray}}\mathbb{E}_{P}\big{[}d_{B}(X,\bar{X})\big{]}\!+\!I_{P}(X;\bar{X})\!-\!R,$ (5)
where $\!d_{B}(x,\bar{x})\!\!\triangleq\!\!-\log\sum_{y}\sqrt{W(y|x)W(y|\bar{x})}$ is the Bhattacharyya distance. This expression is referred to as “weak” since it is looser than the Csiszár-Körner-Marton (CKM) expurgated exponent [9], given by
$E_{\mathrm{ex}}(Q,R)$ $\triangleq\!\!\!\min_{\begin{subarray}{c}P_{X\bar{X}}\\ P_{X}=P_{\bar{X}}=Q\\ I_{P}(X;\bar{X})\leq R\end{subarray}}\!\!\!\mathbb{E}_{P}\!\big{[}d_{B}(X,\bar{X})\big{]}\!+\!I_{P}(X;\bar{X})\!-\!R.$ (6)
The dual expression of (5), as given in [5, Eq. 9], is
$E_{\mathrm{ex}}^{\prime}(Q,R)=\sup_{\rho\geq 1}E_{\mathrm{x}}^{\prime}({Q},\rho)-\rho R,$ (7)
where
$E_{\mathrm{x}}^{\prime}({Q},\rho)\triangleq\min_{Q^{\prime}}-\rho\sum_{x}Q(x)\log\sum_{\bar{x}}Q^{\prime}(\bar{x})\,e^{-\frac{d_{B}(x,\bar{x})}{\rho}}.$ (8)
Using the packing lemma [1, Lemma 6], Csiszár [4] derived two achievable expurgated exponents for joint source-channel coding. The first exponent was derived by assigning the same composition to all source messages, in a similar way to Gallager’s [6, Problem 5.16] for the i.i.d. ensemble. The expression is given by
$E_{J,1}^{\mathrm{ex}}(t,P_{V})\triangleq\max_{Q}\min_{R}\left[te\!\left(\frac{R}{t},P_{V}\right)+E_{\mathrm{ex}}(Q,R)\right].$ (9)
The second exponent requires partitioning of the source sequences and is given by
$E_{J,2}^{\mathrm{ex}}(t,P_{V})\triangleq\min_{R}\left[te\!\left(\frac{R}{t},P_{V}\right)+\max_{Q}E_{\mathrm{ex}}^{\prime}(Q,R)\right].$ (10)
It is not known which of the two exponents is higher. Csiszár partitions the source sequences according to their type (i.e., $m=N_{k}$), assigning a distinct codeword distribution to each one. A more general partitioning was later proposed in [3], where they derived the maximum a posteriori (MAP) random-coding error exponent for arbitrary partitioning of the source sequence set. It was further shown that a specific two-class partitioning suffices to recover Csiszár’s random-coding exponent. In this work, we consider a less general partitioning, where each class contains one or more full source type classes. This setting includes Csiszár’s construction as a special case. We then show that two carefully chosen classes are sufficient to achieve $E_{J,2}^{\mathrm{ex}}(t,P_{V})$.
The work in [5] revisits expurgation in joint source-channel coding and rederives Csiszár’s exponents from [4] using Gallager’s classical expurgation method [6, Sec. 5.7]. The key result from[5] is summarized in the following lemma.
Lemma 1.
There exists a code $\mathcal{C}$ such that for any source type $P_{i}\in\mathcal{P}_{k}(\mathcal{V})$ and any sequence $\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})$, the error probability satisfies
$p_{e}(\boldsymbol{v},{\mathcal{C}})\,\,\dot{\leq}\,\,\sum_{j=1}^{N_{k}}\,\mathbb{E}\Big{[}p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}^{\frac{1}{\rho_{ij}}}\Big{]}^{\rho_{ij}},$ (11)
where $\rho_{ij}>0$, and $p_{e}\big{(}\boldsymbol{v},j,\mathcal{C}\big{)}$ refers to the probability that, given $\boldsymbol{V}=\boldsymbol{v}$, there exists some $\boldsymbol{\bar{v}}\in\mathcal{T}^{k}(P_{j})$ that yields a decoding metric at least as high as that of $\boldsymbol{v}$. Moreover, the expectation is taken with respect to the given (not necessarily distinct) codeword distributions $\{P_{\boldsymbol{X}}^{(j)}\}_{j=1}^{N_{k}}$.
III Main Results
In this section, we present the main results. We begin with the results in the primal domain, followed by direct derivations of their counterparts in the dual domain; their equivalence follows from Lagrange duality [10]. Direct dual-domain derivations have the advantages of remaining valid for arbitrary alphabets, and error exponents remain achievable for any choice of the optimization parameters.
In this paper, we consider the partitioning of the source sequences into $m$ disjoint sets $\mathcal{A}_{1},\ldots,\mathcal{A}_{m}$, each containing one or more full type classes, associated with a codeword distribution $Q_{c}\in\mathcal{P}_{n}(\mathcal{X})$, where the size of the type class must satisfy $\left|\mathcal{T}^{n}(Q_{c})\right|\geq\left|\mathcal{A}_{c}\right|$ for all $c=1,\ldots,m$. We denote by $\Lambda_{c}\subset\mathcal{P}_{k}(\mathcal{V})$ the set of source types included in $\mathcal{A}_{c}$. For each source sequence $\boldsymbol{v}\in\mathcal{A}_{c}$, the codeword $\boldsymbol{x}_{\boldsymbol{v}}\in\mathcal{X}^{n}$ is drawn independently and uniformly from the type class $\mathcal{T}^{n}(Q_{c})$.
III-A Primal-Domain
As a first step, we consider an arbitrary partition of the source sequences and derive an error bound for the code in Lemma 1.
Lemma 2.
Let $\mathcal{A}_{1},...,\mathcal{A}_{m}$ be a given arbitrary partition of the source sequence, each associated with a codeword distribution $Q_{c}\in\mathcal{P}_{n}(\mathcal{X})$. Then there exists a codebook such that
$p_{e}\,\dot{\leq}\sum_{c=1}^{m}\sum_{i\in\Lambda_{c}}\exp\!\left(\!-n\!\left[t{e}\!\left(\frac{R_{i}}{t},P_{V}\!\right)\!+\!E_{\mathrm{ex}}^{\prime}(Q_{c},R_{i})\right]\!\right).$ (12)
In the next theorem, we consider the partitioning that maximizes the exponent in Lemma 2.
Theorem 1.
For a given set of codeword distributions $\mathcal{Q}_{m}=\{Q_{1},Q_{2},\ldots,Q_{m}\}$, there exists a partition of the source sequences into $m$ disjoint sets such that the following exponent is achievable
$\!\!\!{E}_{J,\mathrm{ex}}^{\mathrm{primal}}(t,\mathcal{Q}_{m})$ $\!\triangleq\min_{R}\!\left[t{e}\Big{(}\frac{{R}}{t},P_{V}\Big{)}\!+\!\max_{{Q}\in\mathcal{Q}_{m}}\!\!{E}_{\mathrm{ex}}^{\prime}({Q},{R})\right].$ (13)
Remark 1.
If the distribution in $\mathcal{Q}_{m}$ that maximizes ${E}_{\mathrm{ex}}^{\prime}({Q},{R})$ is independent of $R$, then
$\!\!{E}_{J,\mathrm{ex}}^{\mathrm{primal}}(t,\mathcal{Q}_{m})$ $\!=\!\!\max_{{Q}\in\mathcal{Q}_{m}}\!\min_{R}\!\left[t{e}\Big{(}\frac{{R}}{t},P_{V}\Big{)}\!+\!{E}_{\mathrm{ex}}^{\prime}({Q},{R})\right].$ (14)
Note that the exponent in (14) can be achieved without partitioning, by generating all codewords from the type class of the distribution in $\mathcal{Q}_{m}$ that maximizes ${E}_{\mathrm{ex}}^{\prime}({Q},{R})$. However, if it were known in advance that all codewords share the same composition, then ${E}_{\mathrm{ex}}^{\prime}(Q,R)$ in (14) would be replaced by ${E}_{\mathrm{ex}}(Q,R)$, resulting in a higher exponent.
III-B Dual-Domain
The following lemma presents the analogue of Lemma 2 in the dual domain.
Lemma 3.
Consider a given partitioning of the source sequences as described in Lemma 2. Then, there exists a codebook such that for any choice of parameters $\{\lambda_{ic},\rho_{ic}\}$ satisfying $\lambda_{ic}>0$ and $\rho_{ic}\geq 1$,
$\begin{split}&p_{e}\,\dot{\leq}\,\sum_{c=1}^{m}\sum_{i\in\Lambda_{c}}\exp\!\bigg{(}\!\!-n\Big{[}{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{ic}\big{)}+\big{(}\lambda_{ic}-\rho_{ic}\big{)}{R}_{i}\\ &\hskip 99.58464pt-t{E}_{\mathrm{s}}\big{(}\lambda_{ic},P_{V}\big{)}\Big{]}\bigg{)}.\end{split}$ (15)
The next corollary expresses the result of Lemma 3 in the same form as Theorem 1 in [3].
Corollary 1.
The expression in (15) can also be expressed as
$\!p_{e}\,\dot{\leq}\sum_{c=1}^{m}\exp\!\left(\!-\!\sup_{\rho_{c}\geq 1}\!\left\{n{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{c}\big{)}\!-\!tE_{s}^{(c)}\big{(}\rho_{c},P^{k}\big{)}\right\}\!\right),$ (16)
where
$E_{s}^{(c)}\big{(}\rho,P^{k}\big{)}\triangleq\log\Big{(}\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}P^{k}\big{(}\boldsymbol{v}\big{)}^{\frac{1}{1+\rho}}\Big{)}^{1+\rho}.$ (17)
Next theorem considers the partitioning that maximizes the exponent in Lemma 3.
Theorem 2.
Consider a given set of codeword distributions $\mathcal{Q}_{m}=\{Q_{1},Q_{2},\ldots,Q_{m}\}$, as described in Theorem 1. Then, there exists a partition of the source sequences into $m$ disjoint sets such that the following exponent is achievable
${E}_{J,\mathrm{ex}}^{\mathrm{dual}}(t,\mathcal{Q}_{m})\triangleq\sup_{\lambda\geq 1}\Big{\{}\overline{E}^{\prime}_{\mathrm{x}}\big{(}\mathcal{Q}_{m},\lambda\big{)}-t{E}_{\mathrm{s}}\big{(}\lambda,P_{V}\big{)}\Big{\}},$ (18)
where
$\overline{E}^{\prime}_{\mathrm{x}}\big{(}\mathcal{Q}_{m},\lambda\big{)}\triangleq\min_{R}\sup_{\rho\geq 1}\,E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)+\big{(}\lambda-\rho\big{)}R,$ (19)
and $E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)\triangleq\max_{Q\in\mathcal{Q}_{m}}{E}_{\mathrm{x}}^{\prime}(Q,\rho)$.
The next result extends [2, Lemma 3] to the expurgated exponent for a given set of codeword distributions.
Theorem 3.
The function $\overline{E}_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$ is the concave hull on the interval $[1,\infty)$ of the function $E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$.
Remark 2.
If $E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})$ is concave in $\rho$, or equivalently, if the distribution in $\mathcal{Q}_{m}$ that maximizes $E_{\mathrm{x}}^{\prime}(\rho,Q)$ is independent of $\rho$, then $\overline{E}_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})=E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})$, and hence
$E_{J,\mathrm{ex}}^{\mathrm{dual}}(t,\mathcal{Q}_{m})=\sup_{\rho\geq 1}\Big{\{}E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})-tE_{\mathrm{s}}(\rho,P_{V})\Big{\}}.$ (20)
The next corollary considers Csiszár’s setting, where the partitioning is based on source types, and presents the corresponding dual expression of $E_{J,2}^{\mathrm{ex}}(t,P_{V})$ given in (10).
Corollary 2.
If the partitioning is based on source type classes and the associated distributions $\{Q_{i}\}_{i=1}^{N_{k}}$ are chosen according to
$\!\!\!Q_{i}\!=\underset{\tilde{Q}}{\operatorname{argmax}}\!\!\!\sup_{\rho\geq 1,\lambda>0}\!\!{E}_{\mathrm{x}}^{\prime}\big{(}\tilde{Q},\rho\big{)}\!+\!\big{(}\lambda\!-\!\rho\big{)}R_{i}\!-\!t{E}_{\mathrm{s}}\big{(}\lambda,P_{V}\big{)},$ (21)
then
$E_{J,\mathrm{ex}}^{\mathrm{dual}}\big{(}t,\{Q_{i}\}\big{)}=\sup_{\lambda\geq 1}\Big{\{}\overline{{E}}_{\mathrm{x}}^{\prime}\big{(}\lambda\big{)}-t{E}_{\mathrm{s}}\big{(}\lambda,P_{V}\big{)}\Big{\}},$ (22)
where $\overline{{E}}_{\mathrm{x}}^{\prime}\big{(}\lambda\big{)}$ is the concave hull of ${E}_{\mathrm{x}}^{\prime}\big{(}\lambda\big{)}\triangleq\max_{Q}{E}_{\mathrm{x}}^{\prime}\big{(}Q,\!\lambda\big{)}$.
III-C Two-Class Partition of the Source Sequences
Now, we assume that we only have two classes, i.e. $m=2$.
Lemma 4.
For a pair of distributions $\mathcal{Q}=\{Q,Q^{\prime}\}$, the partition that achieves $E_{J,\mathrm{ex}}^{\mathrm{dual}}(t,\mathcal{Q})$ (or $E_{J,\mathrm{ex}}^{\mathrm{primal}}(t,\mathcal{Q})$), is
$\mathcal{A}_{1}$ =$=\left\{\boldsymbol{v}:\big{|}\mathcal{T}^{k}(\hat{P}_{\boldsymbol{v}})\big{|}\leq e^{nR_{0}}\right\}$ (23)
$\mathcal{A}_{2}$ =$=\left\{\boldsymbol{v}:\big{|}\mathcal{T}^{k}(\hat{P}_{\boldsymbol{v}})\big{|}>e^{nR_{0}}\right\},$ (24)
for some $R_{0}>0$ that depends on $t,P_{V}$, $W$, and $\mathcal{Q}$.
The following corollary is the expurgated analogue of [3, Corollary 1]. As in the random-coding case, a direct application of Carathéodory’s theorem [11, Cor. 17.1.5] implies that any point on the graph of $\overline{E}_{\mathrm{x}}^{\prime}(\lambda)$ can be expressed as a convex combination of two points on the graph of ${E_{\mathrm{x}}^{\prime}}(\lambda)$. As a result, there exists a pair of distributions $\{Q,Q^{\prime}\}$ such that both points lie on the graph of ${E_{\mathrm{x}}^{\prime}}(\lambda,\{Q,Q^{\prime}\})$. Optimizing over all such pairs gives the following result.
Corollary 3.
There exists a partition of the source sequences into two classes, each associated with a distribution from a given pair, such that $E_{J,2}^{\mathrm{ex}}(t,P_{V})$ in (10) is achievable.
An immediate implication of Corollary 3 is that $E_{J,2}^{\mathrm{ex}}(t,P_{V})$ can be achieved with only two classes, in contrast to Csiszár’s construction [4], which requires $m=N_{k}$ classes.
IV Proofs
Due to space constraints, we present only the direct derivations of the dual-domain results, as they are less straightforward and not as immediate as those in the primal domain. The primal-domain case follows by similar arguments, using standard tools from the method of types.
IV-A Proof of Lemma 3
The proof closely follows the approach in [5, Sec. III.C]. We consider the code from Lemma 1, and use the following decoding metric
$q(\boldsymbol{v},\boldsymbol{x},\boldsymbol{y})=W^{n}(\boldsymbol{y}|\boldsymbol{x})\exp\!\big{(}2kD(P_{i}\|P_{V})\big{)}$ (25)
if $\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})$. Then, for any $s>0$, we can upper bound $p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}$ as
$\!\!p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}\!$ $=\mathbb{P}\!\left[\bigcup_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\!\!\!q(\boldsymbol{\bar{v}},\boldsymbol{x}_{\bar{\boldsymbol{v}}},\boldsymbol{Y})\geq q(\boldsymbol{v},\boldsymbol{x}_{\boldsymbol{v}},\boldsymbol{Y})\right]$ (26)
$=\mathbb{P}\!\left[\bigcup_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\!\!\!q(\boldsymbol{\bar{v}},\boldsymbol{x}_{\bar{\boldsymbol{v}}},\boldsymbol{Y})^{s}\geq q(\boldsymbol{v},\boldsymbol{x}_{\boldsymbol{v}},\boldsymbol{Y})^{s}\right]$ (27)
$\leq\sum_{\boldsymbol{y}}W^{n}(\boldsymbol{y}|\boldsymbol{x}_{\boldsymbol{v}})\!\!\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\left(\frac{q(\boldsymbol{\bar{v}},\boldsymbol{x}_{\bar{\boldsymbol{v}}},\boldsymbol{y})}{q(\boldsymbol{v},\boldsymbol{x}_{\boldsymbol{v}},\boldsymbol{y})}\right)^{s}.$ (28)
Substituting the decoding metric and setting $s=\frac{1}{2}$, we obtain
$\begin{split}p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}\,&\leq\left(\frac{\exp\left(kD\left(P_{j}\|P_{V}\right)\right)}{\exp\left(kD\left(P_{i}\|P_{V}\right)\right)}\right)\\ &\times\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\sum_{\boldsymbol{y}}\sqrt{W^{n}(\boldsymbol{y}|\boldsymbol{x}_{\boldsymbol{v}})W^{n}(\boldsymbol{y}|\boldsymbol{x}_{\bar{\boldsymbol{v}}})}.\end{split}$ (29)
If we assume $\rho_{ij}\geq 1$ for all $(i,j)$, we can use Hölder’s inequality [6, Problem 4.15.f] to move the summation over $\boldsymbol{\bar{v}}$ outside, as follows
$\begin{split}\!\!p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}^{\frac{1}{\rho_{ij}}}\!&\leq\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)^{\frac{1}{\rho_{ij}}}\\ &\times\!\!\!\!\!\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}(P_{j})}\prod_{l=1}^{n}\!\left[\sum_{y}\!\sqrt{W(y|x_{\boldsymbol{v}}^{l})W(y|x_{\boldsymbol{\bar{v}}}^{l})}\right]^{\frac{1}{\rho_{ij}}}\!\!\!,\end{split}$ (30)
where $x_{\boldsymbol{v}}^{l}$ denotes the $l$-th element of $\boldsymbol{x}_{\boldsymbol{v}}$. We now take the expectation of the above expression
$\begin{split}\mathbb{E}\left[p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}^{\frac{1}{\rho_{ij}}}\right]\,&\leq\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)^{\frac{1}{\rho_{ij}}}\\ &\times\mathbb{E}\left[\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\prod_{l=1}^{n}\,e^{\frac{-d_{B}(X_{l},\bar{X}_{l})}{\rho_{ij}}}\right],\end{split}$ (31)
where the expectation is taken with respect to $P_{\boldsymbol{X}}^{(\hat{i})}\times P_{\boldsymbol{X}}^{(\hat{j})}$, where $P_{\boldsymbol{X}}^{(c)}$ denotes the uniform distribution over the type class $\mathcal{T}^{n}(Q_{c})$. Here, $\hat{i}$ denotes the index of the class to which source type $i$ belongs. We can evaluate this expectation as follows
$\begin{split}&\mathbb{E}\left[\,\prod_{l=1}^{n}\,e^{\frac{-d_{B}(X_{l},\bar{X}_{l})}{\rho_{ij}}}\right]\\ &\hskip 28.45274pt=\sum_{(\boldsymbol{x},\bar{\boldsymbol{x}})}P_{\boldsymbol{X}}^{(\hat{j})}({\boldsymbol{x}})P_{\boldsymbol{X}}^{(\hat{i})}(\bar{\boldsymbol{x}})\prod_{k=1}^{n}\,e^{\frac{-d_{B}(x_{l},\bar{x}_{l})}{\rho_{ij}}}.\end{split}$ (32)
We use Lemma 2.1 of [12] to rewrite the first expectation as if the distribution were i.i.d. with $Q_{\hat{i}}$. This gives
$\begin{split}&\mathbb{E}\left[\,\prod_{l=1}^{n}\,e^{\frac{-d_{B}(X_{l},\bar{X}_{l})}{\rho_{ij}}}\right]\\ &\hskip 14.22636pt\,\dot{\leq}\sum_{\boldsymbol{x}\in\mathcal{T}^{n}(Q_{\hat{j}})}\frac{1}{\big{|}\mathcal{T}^{n}(Q_{\hat{j}})\big{|}}\prod_{l=1}^{n}\sum_{\bar{x}}Q_{\hat{i}}(x)\,e^{\frac{-d_{B}(x_{l},\bar{x})}{\rho_{ij}}}\end{split}$ (33)
$\hskip 14.22636pt\doteq\,\prod_{x\in\mathcal{X}}\left[\sum_{\bar{x}}Q_{\hat{i}}(x)\,e^{\frac{-d_{B}(x,\bar{x})}{\rho_{ij}}}\right]^{nQ_{\hat{j}}(x)}$ (34)
$\hskip 14.22636pt=\exp\left(n\sum_{x}Q_{\hat{j}}(x)\log\sum_{\bar{x}}Q_{\hat{i}}(x)\,e^{\frac{-d_{B}(x,\bar{x})}{\rho_{ij}}}\right).$ (35)
Note that the bound in (35) does not depend on $\boldsymbol{\bar{v}}$, and so
$\begin{split}&\,\mathbb{E}\left[\sum_{\bar{\boldsymbol{v}}\in\mathcal{T}^{k}(P_{j})}\prod_{k=1}^{n}\,e^{\frac{-d_{B}(X_{k},\bar{X}_{k})}{\rho_{ij}}}\right]\\ &\,\,\dot{\leq}\exp\!\left(-n\!\left[-\sum_{x}Q_{\hat{j}}(x)\log\sum_{\bar{x}}Q_{\hat{i}}(x)\,e^{\frac{-d_{B}(x,\bar{x})}{\rho_{ij}}}\!\!-R_{j}\right]\right).\end{split}$ (36)
Raising both sides to the power $\rho_{ij}$ and taking the worst-case choice of $Q_{\hat{i}}$, we obtain
$\mathbb{E}\left[p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}^{\frac{1}{\rho_{ij}}}\right]^{\rho_{ij}}$
$\leq\!\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)\exp\!\left(\!-n\!\left[E_{\mathrm{x}}^{\prime}({Q_{\hat{j}}},\rho_{ij})\!-\!\rho_{ij}R_{j}\right]\right).$ (37)
Having evaluated the expectation in (11), we now obtain the following bound on the error probability of the code in Lemma 1 as follows
$p_{e}(\mathcal{C})$ $=\sum_{\boldsymbol{v}}P^{k}(\boldsymbol{v})\,p_{e}(\boldsymbol{v},\mathcal{C})$ (38)
˙$\begin{split}&\,\,\dot{\leq}\,\,\sum_{(i,j)}\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}P^{k}(\boldsymbol{v})\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)\\ &\hskip 28.45274pt\times\exp\left(-n\left[E_{\mathrm{x}}^{\prime}({Q_{\hat{j}}},\rho_{ij})-\rho_{ij}R_{j}\right]\right).\end{split}$ (39)
Observe that $\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}P^{k}(\boldsymbol{v})\doteq\exp(kD(P_{i}\|P_{V}))$. This implies that
$\!\!\!\!\!\!\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}\!\!\!\!\!P^{k}(\boldsymbol{v})\!\left(\frac{\exp(kD(P_{j}\|P_{V}))}{\exp(kD(P_{i}\|P_{V}))}\right)$ $\!\doteq\exp(kD(P_{j}\|P_{V})).$ (40)
We can further simplify this for any $\lambda_{j}>0$ as follows
$\begin{split}&\exp\big{(}kD(P_{j}\|P_{V})\big{)}\\ &\hskip 7.11317pt=\exp\left(k\sum_{v}P_{j}(v)\log\frac{P_{V}(v)}{P_{j}(v)}\right)\end{split}$ (41)
$\hskip 7.11317pt=\exp\!\left(k(1+\lambda_{j})\sum_{v}P_{j}(v)\log\frac{P_{V}(v)^{\frac{1}{1+\lambda_{j}}}}{P_{j}(v)P_{j}(v)^{\frac{-\lambda_{j}}{1+\lambda_{j}}}}\right).$ (42)
The last expression can be written as
$\begin{split}&\exp\left(k(1\!+\!\lambda_{j})\sum_{v}P_{j}(v)\log\frac{P_{V}(v)^{\frac{1}{1+\lambda_{j}}}}{P_{j}(v)P_{j}(v)^{\frac{-\lambda_{j}}{1+\lambda_{j}}}}\right)\\ &=\exp\!\left(\!n\!\left[t(1\!+\!\lambda_{j})\!\sum_{v}P_{j}(v)\log\frac{P_{V}(v)^{\frac{1}{1+\lambda_{j}}}}{P_{j}(v)}\!-\!\lambda_{j}R_{j}\right]\right)\end{split}$ (43)
$\leq\exp\!\left(\!n\!\left[t(1\!+\!\lambda_{j})\!\log\sum_{v}P_{j}(v)\frac{P_{V}(v)^{\frac{1}{1+\lambda_{j}}}}{P_{j}(v)}\!-\!\lambda_{j}R_{j}\right]\right)$ (44)
$=\exp\left(-n\Big{[}\lambda_{j}R_{j}\!-\!tE_{s}(\lambda_{j},P_{V})\Big{]}\right),$ (45)
where we applied Jensen’s inequality in (44). By setting $\rho_{ij}=\rho_{j}$, we can upper bound the expression in (39) as
$\begin{split}&p_{e}(\mathcal{C})\,\,\dot{\leq}\,\,\sum_{j=1}^{N_{k}}\exp\bigg{(}-n\Big{[}{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{\hat{j}},\rho_{j}\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda_{j}-\rho_{j}\big{)}{R}_{j}-t\mathrm{E}_{\mathrm{s}}\big{(}\lambda_{j},P_{V}\big{)}\Big{]}\bigg{)}\end{split}$ (46)
$\begin{split}&\hskip 28.45274pt=\sum_{c=1}^{m}\sum_{i\in\Lambda_{c}}\exp\bigg{(}-n\Big{[}{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{ic}\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda_{ic}-\rho_{ic}\big{)}{R}_{i}-t\mathrm{E}_{\mathrm{s}}\big{(}\lambda_{ic},P_{V}\big{)}\Big{]}\bigg{)}.\end{split}$ (47)
To prove Corollary 1, observe that by applying (40), the bound in (39) can be rewritten as
$p_{e}(\mathcal{C})$$\,\dot{\leq}\,\sum_{i=1}^{N_{k}}\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}\frac{P^{k}(\boldsymbol{v})}{\hat{P}_{\boldsymbol{v}}(\boldsymbol{v})^{\rho_{i}}}\exp\big{(}-n\,{E}_{\mathrm{x}}^{\prime}(Q_{\hat{i}},\rho_{i})\big{)}$ (48)
$=\sum_{c=1}^{m}\left(\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}\frac{P^{k}(\boldsymbol{v})}{\hat{P}_{\boldsymbol{v}}(\boldsymbol{v})^{\rho_{c}}}\right)\exp\!\big{(}\!-\!n\,{E}_{\mathrm{x}}^{\prime}(Q_{c},\rho_{c})\big{)}.$ (49)
The inner sum can be upper bounded as
$\!\!\!\!\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}\frac{P^{k}(\boldsymbol{v})}{\hat{P}_{\boldsymbol{v}}(\boldsymbol{v})^{\rho}}$ $=\sum_{i\in\Lambda_{c}}\sum_{\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})}\frac{P^{k}(\boldsymbol{v})}{P_{i}^{k}(\boldsymbol{v})^{\rho}}$ (50)
$\doteq\sum_{i\in\Lambda_{c}}\exp\left(k\sum_{v}P_{i}(v)\log\frac{P_{V}(v)}{P_{i}(v)^{1+\rho}}\right)$ (51)
$=\sum_{i\in\Lambda_{c}}\!\left[\exp\!\left(\!k\!\sum_{v}\!P_{i}(v)\!\log\!\frac{P_{V}(v)^{\frac{1}{1+\rho}}}{P_{i}(v)}\!\right)\!\right]^{1+\rho}$ (52)
$=\sum_{i\in\Lambda_{c}}\left[\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}P^{k}(\boldsymbol{v})^{\frac{1}{1+\rho}}\right]^{1+\rho}$ (53)
$\leq\left[\sum_{i\in\Lambda_{c}}\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}P^{k}(\boldsymbol{v})^{\frac{1}{1+\rho}}\right]^{1+\rho}$ (54)
$=\left[\sum_{\boldsymbol{v}\in\mathcal{A}_{c}}P^{k}(\boldsymbol{v})^{\frac{1}{1+\rho}}\right]^{1+\rho},$ (55)
where in (51), we used the fact that $|\mathcal{T}^{k}(P_{i})|\doteq\exp(kH(P_{i}))$, and in (54), we applied the inequality $\sum_{i}a_{i}^{r}\leq(\sum_{i}a_{i})^{r}$ for $r\geq 1$ [6, Problem 4.15.f].
IV-B Proof of Theorem 2
We first set $\rho_{ic}=\rho_{c}$ and $\lambda_{ic}=\lambda_{c}$. To find the optimal partitioning, assume each class is associated with given values of $\rho_{c}$ and $\lambda_{c}$, which will be optimized later. The source type class $\mathcal{T}^{k}(P_{i})$ is then assigned to class $c^{\star}$ if
$\!c^{\star}\!=\!\underset{c\in\{1,\ldots,m\}}{\operatorname{argmax}}E_{\mathrm{x}}^{\prime}(\mathrm{Q}_{c},\rho_{c})\!+\!(\lambda_{c}\!-\!\rho_{c}){R}_{i}-t{E}_{\mathrm{s}}(\lambda_{c},P_{V}).$ (56)
This means that
$\begin{split}&p_{e}(\mathcal{C})\,\,\dot{\leq}\,\,\sum_{j=1}^{N_{k}}\exp\bigg{(}-n\bigg{[}\max_{c}\,{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{c}\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda_{c}-\rho_{c}\big{)}{R}_{j}-t\mathrm{E}_{\mathrm{s}}\big{(}\lambda_{c},P_{V}\big{)}\bigg{]}\bigg{)}\end{split}$ (57)
˙$\begin{split}&\hskip 24.18501pt\,\,\dot{\leq}\,\exp\bigg{(}-n\bigg{[}\min_{R}\max_{c}\,{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{c}\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda_{c}-\rho_{c}\big{)}{R}-t\mathrm{E}_{\mathrm{s}}\big{(}\lambda_{c},P_{V}\big{)}\bigg{]}\bigg{)}.\end{split}$ (58)
Optimizing over the parameters yields
$\begin{split}&p_{e}(\mathcal{C})\,\dot{\leq}\,\exp\bigg{(}-n\bigg{[}\min_{R}\max_{\lambda>0}\max_{\rho\geq 1}\max_{Q\in\mathcal{Q}_{m}}\,{E}_{\mathrm{x}}^{\prime}\big{(}{Q},\rho\big{)}\\ &\hskip 56.9055pt+\big{(}\lambda-\rho\big{)}{R}-t{E}_{\mathrm{s}}\big{(}\lambda,P_{V}\big{)}\bigg{]}\bigg{)}.\end{split}$ (59)
Since $E_{s}(\lambda,P_{V})$ is convex in $\lambda$, we can apply Sion’s minimax theorem [13] to exchange the order of $\min_{R}$ and $\max_{\lambda}$, which gives
$\begin{split}&p_{e}(\mathcal{C})\,\dot{\leq}\,\exp\bigg{(}-n\bigg{[}\max_{\lambda>0}\min_{R}\max_{\rho\geq 1}\max_{Q\in\mathcal{Q}_{m}}\,{E}_{\mathrm{x}}^{\prime}\big{(}{Q},\rho\big{)}\\ &\hskip 71.13188pt+\big{(}\lambda-\rho\big{)}{R}-t{E}_{\mathrm{s}}\big{(}\lambda,P_{V}\big{)}\bigg{]}\bigg{)}.\end{split}$ (60)
Observe that if $\lambda<1$, then $(\lambda-\rho)R<0$, which implies that ${R}\rightarrow\infty$. Hence, the range of $\lambda$ is restricted to $\lambda\geq 1$.
IV-C Proof of Theorem 3
Let $g(\lambda)\triangleq E_{\mathrm{x}}^{\prime}(\lambda,\mathcal{Q}_{m})$. Adopting from [2], the concave conjugate of $g(\cdot)$ over $[1,\infty)$ is given by [11]
$g^{\star}(\lambda)=\inf_{\rho\geq 1}\Big{\{}\lambda\rho-E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})\Big{\}}.$ (61)
The concave hull of $g(\cdot)$ is given by its biconjugate $g^{\star\star}(\lambda)$. Hence, we obtain
$g^{\star\star}(\lambda)$ $=\inf_{R}\bigg{\{}\lambda R-\inf_{\rho\geq 1}\Big{\{}\rho R-E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})\Big{\}}\bigg{\}}$ (62)
$=\inf_{R}\sup_{\rho\geq 1}\Big{\{}E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})+(\lambda-\rho)R\Big{\}}.$ (63)
This argument closely follows the proof of [2, Lemma 3].
IV-D Proof of Lemma 4
Assume the two classes are associated with given parameters $\lambda_{c}$ and $\rho_{c}$. If $\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})$ is assigned to class 1, then by (56), we have
$\begin{split}&E_{\mathrm{x}}^{\prime}({Q}_{1},\rho_{1})+(\lambda_{1}-\rho_{1}){R}_{i}-t{E}_{\mathrm{s}}(\lambda_{1},P_{V})\\ &\hskip 21.33955pt\geq E_{\mathrm{x}}^{\prime}({Q}_{2},\rho_{2})+(\lambda_{2}-\rho_{2}){R}_{i}-t{E}_{\mathrm{s}}(\lambda_{2},P_{V}).\end{split}$ (64)
Rearranging terms, this implies
$\!R_{i}\!\leq\!\frac{E_{\mathrm{x}}^{\prime}({Q}_{1},\rho_{1})\!-\!E_{\mathrm{x}}^{\prime}({Q}_{2},\rho_{2})\!+\!t\left[{E}_{\mathrm{s}}(\lambda_{2},P_{V})\!-\!t{E}_{\mathrm{s}}(\lambda_{1},P_{V})\right]}{\big{(}\lambda_{2}-\rho_{2}\big{)}-\big{(}\lambda_{1}-\rho_{1}\big{)}}.$ (65)
Similarly, if $\boldsymbol{v}\in\mathcal{T}^{k}(P_{i})$ is assigned to class 2, the inequality in (65) is reversed.